# Generative Data Mining with Longtail-Guided Diffusion

### v1

@@ Abstract
It is difficult to anticipate the myriad challenges that a predictive model will encounter once deployed. Common practice entails a reactive, cyclical approach: model deployment, data mining, and retraining. We instead develop a proactive longtail discovery process by imagining additional data during training. In particular, we develop general model-based longtail signals, including a differentiable, single forward pass formulation of epistemic uncertainty that does not impact model parameters or predictive performance but can flag rare or hard inputs. We leverage these signals as guidance to generate additional training data from a latent diffusion model in a process we call Longtail Guidance (LTG). Crucially, we can perform LTG without retraining the diffusion model or the predictive model, and we do not need to expose the predictive model to intermediate diffusion states. Data generated by LTG exhibit semantically meaningful variation, yield significant generalization improvements on image classification benchmarks, and can be analyzed to proactively discover, explain, and address conceptual gaps in a predictive model.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Diffusion with Longtail Guidance (LTG) generates synthetic data that are difficult or rare for an existing predictive model. The predictive model can then a) be fine-tuned on this data to improve generalization performance, or b) the synthetic data can be analyzed to understand conceptual gaps in the predictive model. Guided generations exhibit more extreme views and lighting compared to unguided generations.
Figure 2: Predictive model performance (left axis, red) and model-based longtail signals $f_{\phi}^{\text{lt}}$ (right axis, green) on synthetic data generated by varying Longtail Guidance weights for two types of longtail signals: Epistemic (left plot) and Entropy (right plot). In all cases, guided synthetic data generations exhibit lower correct class probability, lower accuracy, and higher longtail signals compared to unguided generations (guidance weight 0).
@@
We wish to trade additional offline compute for performance by generatively mining additional training data that is hard or rare from the perspective of an existing predictive model. Distinct from adversarial training and robustness approaches (Bansal & Grover, 2023; Zhang et al., 2022a; Gowal et al., 2021; Hendrycks et al., 2021) which imbue models with invariance to distribution shift or small, often imperceptible changes, we wish to generate semantically meaningful, in-distribution examples that cause high model uncertainty.
Foundation generative models such as Stable Diffusion (Rombach et al., 2022), Pixart (Chen et al., 2025), GPT (Brown et al., 2020; OpenAI, 2023), and Claude (Anthropic, 2024) are exposed to Internet-scale data, but it is not obvious how to use them to generate data that is specifically challenging to an existing predictive model. We seek to explicitly couple the vast knowledge contained in foundation generative models with the specific challenges faced by an existing predictive model. But how can we ensure that the generative model produces data that is relevant (has meaningful learning signal) to a given predictive model? Existing approaches, like prompt tuning and reasoning in an embedding space (such as CLIP) (Zhang et al., 2023b; Du et al., 2024) can generate synthetic data that improves model generalization performance up to a point, but they do not condition on the deployed model.
@@
    Longtail Introspection: through VLM analysis of LTG-generated data, we find keywords describing what an existing predictive model struggles with. Keywords can be used to prompt diffusion models for high-value data or to inform future data collection.
@@ 2.1 Epistemic Head
In Figure 3 (left), we introduce a lightweight ensembling technique called the Epistemic Head, which provides a superior longtail signal with no impact on model performance and negligible changes to model parameter count, training time, and inference time (see Supplement A.6). Inspired by LoRA (Hu et al., 2021) and the Hydra architecture (Tran et al., 2020), we duplicate the head of an existing prediction model $K$ times and jointly train with the same loss as the base model but with diversity encouraged through an oracle loss (Guzman-Rivera et al., 2012) that only propagates per-example loss through the best-performing head. Head outputs act as fixed-point samples from the posterior predictive (Eqn 3) under a prior determined by weight initialization and permit differentiable computation of longtail signals in a single forward pass, including $E(y,\Phi)$,
@@
Base model performance is protected by a stop-grad at training time so that, similar to LoRA (Hu et al., 2021), the Epistemic Head does not impact existing model weights. In Figure 5, we demonstrate that longtail signals from the Epistemic Head more effectively indicate rare or hard examples than do entropy or energy. In particular, we train ViT-B ImageNet-LT classifiers according to (Xu et al., 2023) and compare how well our proposed longtail signals (entropy, energy, epistemic) detect rare or hard test examples. Examples are defined as rare if they come from a longtail class. Examples are defined as hard if the model incorrectly predicts the label. In Supplement A.7, we further show that detected examples are disproportionately hard.
@@ 3 Longtail Guidance
Figure 6: Synthetic data generated for 100 ImageNet classes with increasing Longtail Guidance strength, guided by a SOTA ViT ImageNet-LT classifier $f_{\phi}$. Views frequently become more extreme, occluded, or cut off. They also become more difficult for $f_{\phi}$ (see Figure 2)
@@
However, they differ in the learned reverse process: DDPM models it as a Markov chain,
@@
Input : Latent diffusion model $\epsilon_{\theta}(z_{t},t)$, predictor $f_{\phi}(x)$ decoder $D$, noise schedule $\sigma_{1:T}$, weight $w$
@@
It is unintuitive that LTG would work since production model $f_{\phi}$ has only ever trained on clean (data, label) pairs, not intermediate diffusion states. Empirically, however, we find that we can generate synthetic data for which production model $f_{\phi}$ exhibits lower probability of the correct class, lower accuracy, and higher longtail signal. And so long as we do not adjust the longtail guidance weight $w$ too high, the diffusion model reliably generates data that adheres to the expected class label. Figures 16 and Supplement A.10 (high-res) show example longtail synthetic ImageNet data using a SOTA ViT model (Xu et al., 2023) for guidance. Figure 2 quantitatively analyzes model performance and longtail signals on LTG-guided synthetic data generation over multiple runs of $24k$ generations spanning all ImageNet classes. Notably, the probability of the expected class can be decreased to one-half or even one-third of its original value (from the perspective of production model $f_{\phi}$), while staying in-distribution (as evidenced by generalization improvements in Section 4) and the model longtail signal can be more than doubled over baseline (unguided) synthetic data generation. Qualitatively, LTG data exhibit semantically meaningful changes, including more extreme, occluded, or cut off views.
Longtail Guidance is similar to Universal Guidance (Bansal et al., 2023) in that both perform differentiable decoding of latent states (see also (Jiang et al., 2023)). However, Universal Guidance is significantly more expensive and fails to remain in-distribution when guided by model-based longtail signals. In particular, Universal Guidance performs so-called recurrent and backward sampling passes, where recurrent sampling permits each diffusion step to refine its denoising estimate (each time computing a costly gradient) and each backward pass optimizes an objective (within each denoising and recurrent iterate, itself requiring multiple optimization iterates). See Figure 7 for direct comparison.
Although Longtail Guidance is not the first approach to use the predicted, decoded terminal state $D(P(x_{t}))$, we uniquely (to our knowledge) show in Supplement A.8 why the predictive model performing guidance does not need to be trained on intermediate diffusion states: $D(P(x_{t}))$ is closer in distribution to real training data than is $x_{t}$.
LTG requires no training for either the diffusion or the production model. Each can be used off-the-shelf. It’s primary limitation is that the gradient $d\hat{x}_{t}^{0}/d\hat{z}_{t}^{0}$ calculated by decoding is expensive (see Supplement A.6).
@@ 4 Experiments
In Section 4.1, we compare predictive model generalization improvements when training data is augmented with synthetic data generated by Longtail Guidance, when training data is augmented with existing synthetic data generation approaches, and when training data is augmented with traditional data augmentations approaches. In Section 4.2, we reduce synthetic data generated with Longtail Guidance to a set of text descriptions that describe attributes of a predictive model’s longtail. We demonstrate that these descriptions are meaningful by showing that they produce higher-value synthetic data than manually prompt-tuned diffusion. In Supplement A.1, we additionally demonstrate that LTG improves the ImageNet-LT generalization performance of SOTA ViT models that compensate for class-imbalance. In Section 4.3, we examine why Longtail Guidance outperforms existing synthetic data generation approaches.
@@ 4.1 LTG Improves Predictive Model Generalization
We evaluate on seven natural image datasets spanning fine-grained, coarse-grained, and mixed coarse/fine classification tasks. Datasets include ImageNet, ImageNet-V2, ImageNet-A, Stanford Cars (Krause et al., 2013), Oxford Flowers (Nilsback & Zisserman, 2008), Oxford Pets (Parkhi et al., 2012), and Caltech101 (Fei-Fei et al., 2004). We reproduce baseline predictive models (Original) that are not exposed to synthetic data, with comparable performance (within $1\%$) of the baselines used in (Du et al., 2024; Zhang et al., 2023b). [^1] Baseline model checkpoints were not available so we reproduce them using the reported architecture and training recipe. We iteratively perform additional fine-tuning with synthetic training data that is equal in quantity across all conditions, use the same augmentations (random rotation, reflection, and center cropping), and report generalization performance as top-1 accuracy. Results are averaged over three runs. Additional experiment details are in Supplement A.2 and dataset details are in Supplement A.4. We fix hyperparameters (LTG weight, number of Epistemic Heads) across all datasets according to Supplement A.9.
@@
In Table 1, we compare to baselines under the Dataset Expansion task as defined in GIF (Zhang et al., 2023b). For parity, we fine-tune ResNet-50 models for 100 epochs and generate synthetic data equivalent to 20×20$\times$20 × the original dataset size for Caltech, Cars and Flowers, and 30×30$\times$30 × the original dataset size for Pets. GIF generates all synthetic data at once whereas we evenly distribute data generation throughout training epochs to better capture evolving model-based longtail challenges, see Supplement A.2 for details. In each expansion, synthetic examples per class are equal in quantity to the original number of real examples.
In Table 2, we compare to baselines under the in-distribution synthetic data generation task as defined in Dream the Impossible (Du et al., 2024) and discussed in Supplement A.2. For parity, we train ResNet-34 models for 100 epochs and generate 1k synthetic data per class. Models are trained on ImageNet-100 and evaluated on ImageNet-100, ImageNet-A, and ImageNet-V2 subsets defined in Supplement A.4.
@@
In each benchmark, Longtail Guidance provides substantial improvements over existing synthetic data generation methods and data augmentation baselines. In particular, it provides an average $5.6$ points of additional top-1 accuracy compared to GIF or Dream-ID, and an average $15.0$ points of additional top-1 accuracy compared to text-prompted Stable Diffusion. Aggregate improvements from the Epistemic Head over alternative model-based longtail signals (energy, entropy) are mild but significant and come at negligible inference or training time cost as described in Supplement A.6. In Supplement A.5, we show that predictive model generalization can be further improved by engaging in more cycles of fine-tuning and synthetic data generation.
@@
4.2 LTG Produces Meaningful Longtail Data
Refined prompts are used to generate additional synthetic data $x_{\text{inspect}}$ by diffusion (without LTG). We train $f_{\phi}$ on synthetic data generated by Longtail Introspection and compare it to synthetic data generated by manual prompt tuning (40 unique prompts per class, as in (Zhang et al., 2023b); details in Supplement A.3). Table 3 shows that Longtail Introspection significantly outperforms manually prompt-tuned diffusion (defined in Supplement A.2), quantitatively supporting that data generated by LTG exhibit meaningful variation. Figure 8 shows example longtail keywords.
Figure 8: Longtail keywords from LTG-generated Flowers data.
Condition Accuracy
No Synthetic Data 74.1
Manual Prompt Tuning 78.8
Longtail Introspection 84.3
4.3 Discussion
5 Related Work
Diffusion has seen wide success, particularly in image generation, where it outperform GANs in image quality and diversity without suffering from unstable training or mode collapse (Dhariwal & Nichol, 2021). Recent work has seen progress in handling large data dimensions with latent spaces (Chen et al., 2025; Podell et al., 2023; Rombach et al., 2022) or hourglass networks (Crowson et al., 2024), improved sampling (Lu et al., 2022; Karras et al., 2022; Ho et al., 2020; Song et al., 2020a), additional data domains (Ran et al., 2024; Pronovost et al., 2023; Zhong et al., 2023), and personalization (Ruiz et al., 2023; Kumari et al., 2023). Much work has also been done on new forms of guidance (Wallace et al., 2023b; Yu et al., 2023a; Wallace et al., 2023a; Zhang et al., 2023a) beyond just classifier guidance and classifier-free guidance (Dhariwal & Nichol, 2021; Ho & Salimans, 2022). Universal Guidance (Bansal et al., 2023) is most relevant, and is discussed in Section 3.
Model signals have been used for out-of-distribution or adversarial detection (Huang et al., 2021; Hsu et al., 2020; Cohen et al., 2020), particularly model uncertainty (Van Amersfoort et al., 2020) or feature density (Lee et al., 2018). Epistemic uncertainty has been developed in (expensive) Bayesian (Gal & Ghahramani, 2016) or model ensemble contexts (Liu et al., 2019a; Depeweg et al., 2018; Wilson & Izmailov, 2020; Lakshminarayanan et al., 2017), though, to our knowledge, not explicitly in a single, differentiable forward pass as we have done in Section 2.1. Work on longtail robustness has primarily focused on addressing a-priori known class imbalance. Datasets include ImageNet-LT, Places-LT, (Liu et al., 2019b) and iNaturalist (Van Horn et al., 2018). Mitigations include pretraining (He et al., 2022; Bao et al., 2021), distillation (Xiang et al., 2020), reweighted losses (Xu et al., 2023; Ross & Dollár, 2017), or selective (real) data mining (Jiang et al., 2022).
Conclusion
We developed model-based longtail signals that do not impact model weights or performance and are leveraged by Longtail Guidance, a synthetic data generation approach that explicitly conditions on an existing predictive model to generate examples that are rare or hard from that model’s perspective. We show that training on LTG-generated data provides stronger generalization improvements than synthetic data generation and data augmentation baselines over eight datasets. We further demonstrate that longtail synthetic data generations can be rendered into meaningful text descriptions that can aid future (real or synthetic) data collection priorities.
Predictive models are being deployed more than ever before. Increasingly, they will encounter longtail scenarios that human operators cannot predict in advance. Foundation models can be used to mitigate some risk, but we cannot shoehorn the entirety of Internet-scale knowledge into every deployed model due to capacity and compute constraints. By letting an existing predictive model speak for itself, we can expend offline compute to generatively mine high-value synthetic training data. Further reducing these synthetic examples to human- and machine-readable text suggests a future where we can move away from slow, reactive longtail mitigation towards fast, proactive longtail discovery.
Appendix A Appendix

### v2

@@ Abstract
It is difficult to anticipate the myriad challenges that a predictive model will encounter once deployed. Common practice entails a reactive, cyclical approach: model deployment, data mining, and retraining. We instead develop a proactive longtail discovery process by imagining additional data during training. In particular, we develop general model-based longtail signals, including a differentiable, single forward pass formulation of epistemic uncertainty that does not impact model parameters or predictive performance but can flag rare or hard inputs. We leverage these signals as guidance to generate additional training data from a latent diffusion model in a process we call Longtail Guidance (LTG). Crucially, we can perform LTG without retraining the diffusion model or the predictive model, and we do not need to expose the predictive model to intermediate diffusion states. Data generated by LTG exhibit semantically meaningful variation, yield significant generalization improvements on numerous image classification benchmarks, and can be analyzed by a VLM to proactively discover, textually explain, and address conceptual gaps in a deployed predictive model.
Synthetic Data, Longtail, Long Tail, Foundation Model, Diffusion, Guidance, VLM, CLIP, Embedding, Text, Robustness, Uncertainty, Epistemic, Aleatoric, Text, Autolabel, Imagine, Imagination, Dream
@@ 1 Introduction
Figure 1: Diffusion with Longtail Guidance (LTG) generates synthetic data that are difficult or rare for an existing predictive model. The predictive model can be fine-tuned on this data to improve generalization performance and the the synthetic data can be analyzed to understand conceptual gaps in the predictive model. Guided generations exhibit more extreme views compared to unguided generations. Results in Section 4 use low-to-mid guidance weights.
Figure 2: Predictive model performance (left axes, red) and model-based longtail signals $f_{\phi}^{\text{lt}}$ (right axes, green) on synthetic data generated by varying Longtail Guidance weights for model-based longtail signals: Epistemic (left plot) and Entropy (right plot). Guided generations predictably exhibit lower correct class probability, P(Class), lower accuracy, and higher longtail signals, $f_{\phi}^{\text{lt}}$, compared to unguided generations (zero guidance weight), indicating they are more difficult and more longtail from the predictive model’s perspective. We ensure that longtail-guided synthetic data generations do not stray out-of-distribution by ensuring P(Class) is lower than unguided generations but well above zero probability.
@@
We wish to trade additional offline compute for performance by generatively mining additional training data that is hard or rare from the perspective of an existing predictive model. Distinct from adversarial training and robustness approaches (Bansal & Grover, 2023; Zhang et al., 2022a; Gowal et al., 2021; Hendrycks et al., 2021) which imbue models with invariance to distribution shift or small, often imperceptible changes, we aim to generate semantically meaningful, in-distribution examples that cause high model uncertainty.
Foundation generative models such as Stable Diffusion (Rombach et al., 2022), Pixart (Chen et al., 2025), GPT (Brown et al., 2020; OpenAI, 2023), and Claude (Anthropic, 2024) are exposed to Internet-scale data, but it is not obvious how to use them to generate data that is specifically challenging to an existing predictive model. We seek to explicitly couple the vast knowledge contained in foundation generative models with the specific challenges faced by an existing predictive model. But how can we ensure that the generative model produces data that is relevant (has meaningful learning signal) to a given predictive model? Existing approaches, like prompt tuning and reasoning in an embedding space (such as CLIP) (Zhang et al., 2023b; Du et al., 2024) can generate synthetic data that improves model generalization performance up to a point, but they do not condition on the deployed predictive model.
@@
    Longtail Introspection: through VLM analysis of LTG-generated data, we find keywords describing what an existing predictive model struggles with. Keywords can be used to prompt diffusion models for high-value data or to inform future real data collection.
@@ 2.1 Epistemic Head
In Figure 3 (left), we introduce a lightweight ensembling technique called the Epistemic Head, which provides a superior longtail signal with no impact on model performance and negligible changes to model parameter count, training time, and inference time (see Supplement A.5). Inspired by LoRA (Hu et al., 2021) and the Hydra architecture (Tran et al., 2020), we duplicate the head of an existing prediction model $K$ times and jointly train with the same loss as the base model but with diversity encouraged through an oracle loss (Guzman-Rivera et al., 2012) that only propagates per-example loss through the best-performing head. Head outputs act as fixed-point samples from the posterior predictive (Eqn 3) under a prior determined by weight initialization and permit differentiable computation of longtail signals in a single forward pass, including $E(y,\Phi)$,
@@
Base model performance is protected by a stop-grad at training time so that, similar to LoRA (Hu et al., 2021), the Epistemic Head does not impact existing model weights. In Figure 5, we demonstrate that longtail signals from the Epistemic Head more effectively indicate rare or hard examples than do entropy or energy. In particular, we train ViT-B ImageNet-LT classifiers according to (Xu et al., 2023) and compare how well our proposed longtail signals (entropy, energy, epistemic) detect rare or hard test examples. Examples are defined as rare if they come from a longtail class. Examples are defined as hard if the model incorrectly predicts the label. In Supplement A.6, we further show that detected examples are disproportionately hard.
@@ 3 Longtail Guidance
Figure 6: Synthetic data generated for 100 ImageNet classes with increasing Longtail Guidance strength, guided by a SOTA ViT ImageNet-LT classifier $f_{\phi}$. Views frequently become more extreme, occluded, or cut off. They also become more difficult for $f_{\phi}$ (see Figure 2). Best experiment results use low-to-mid Longtail Guidance weights (see Section 4 for details).
@@
However, they differ in the learned reverse process. DDPM models it as a Markov chain,
@@
Input: Latent diffusion model $\epsilon_{\theta}(z_{t},t)$, predictor $f_{\phi}(x)$ latent decoder $D$, noise schedule $\sigma_{1:T}$, weight $w$
@@
It is unintuitive that LTG would work since production model $f_{\phi}$ has only ever trained on clean (data, label) pairs, not intermediate diffusion states. Empirically, however, we find that we can generate synthetic data for which production model $f_{\phi}$ exhibits lower probability of the correct class, lower accuracy, and higher longtail signal. And so long as we do not adjust the longtail guidance weight $w$ too high, the diffusion model reliably generates data that adheres to the expected class label.
Figures 16 and Supplement A.9 show example longtail synthetic ImageNet data using a SOTA ViT classifier (Xu et al., 2023) for guidance. Figure 2 quantitatively analyzes model performance and longtail signals on LTG-guided synthetic data generation over multiple runs of $24k$ generations spanning all ImageNet classes. Notably, the probability of the expected class can be decreased to one-third of its original value (from the perspective of production model $f_{\phi}$), while staying in-distribution (as evidenced by generalization improvements in Section 4, see also Section 3.1) and the model longtail signal can be more than doubled over baseline (unguided) synthetic data generation. Qualitatively, LTG data exhibit semantically meaningful changes, including more extreme, occluded, or cut-off views.
Longtail Guidance is similar to Universal Guidance (Bansal et al., 2023) in that both perform differentiable decoding of latent states (see also (Jiang et al., 2023; Song et al., 2023; Dou & Song, 2024)). However, Universal Guidance is significantly more expensive and fails to remain in-distribution when guided by model-based longtail signals. In particular, Universal Guidance performs so-called recurrent and backward sampling passes, where recurrent sampling permits each diffusion step to refine its denoising estimate (each time computing a costly gradient) and each backward pass optimizes an objective (within each denoising and recurrent iterate, itself requiring multiple optimization iterates). See Figure 7 for direct comparison.
Although Longtail Guidance is not the first approach to use the predicted, decoded terminal state $D(P(x_{t}))$, we uniquely (to our knowledge) show in Supplement A.7 why the predictive model performing guidance does not need to be trained on intermediate diffusion states: $D(P(x_{t}))$ is closer in distribution to real training data than is $x_{t}$.
LTG requires no training for either the diffusion or the production model. Each can be used off-the-shelf. It’s primary limitation is that the gradient $d\hat{x}_{t}^{0}/d\hat{z}_{t}^{0}$ calculated by decoding is expensive (see Supplement A.5).
@@
3.1 Remaining In-Distribution
We ensure that LTG-generated data remain in-distribution but longtail by choosing a guidance weight such that (1) the probability of the desired class is lower than baseline unguided synthetic data generation but nonzero, and (2) that the model-based longtail signal $f_{\phi}^{\text{lt}}$ evaluates higher than it does for baseline unguided synthetic data. Figure 2 shows that both are achievable; in fact, we find that a single weight can be chosen one time and used for all future generations (across classes, datasets and fine-tuning epochs). With this approach, we find that FID (Heusel et al., 2017) scores are lower and generative precision+recall (Kynkäänniemi et al., 2019) are higher when comparing LTG-generated synthetic data to real data (within-classes) than when comparing real data to itself (between classes).
@@ 4 Experiments
In Section 4.1, we compare predictive model generalization improvements when training data is augmented with synthetic data generated by Longtail Guidance, when training data is augmented with existing synthetic data generation approaches, and when training data is augmented with traditional data augmentations approaches. In Section 4.2, we demonstrate that LTG improves generalization performance of SOTA ViT models that compensate for class-imbalance, particularly for non-synthetic longtail data. In Section 4.3, we reduce synthetic data generated with Longtail Guidance to a set of text descriptions that describe attributes of a predictive model’s longtail. We demonstrate that these descriptions are meaningful by showing that they produce higher-value synthetic data than manually prompt-tuned diffusion (in terms of predictive model generalization performance). In Section 4.4, we examine why Longtail Guidance outperforms existing synthetic data generation approaches.
@@ 4.1 LTG Improves Predictive Model Generalization
We evaluate on seven natural image datasets spanning fine-grained, coarse-grained, and mixed coarse/fine classification tasks. Datasets include ImageNet, ImageNet-V2, ImageNet-A, Stanford Cars (Krause et al., 2013), Oxford Flowers (Nilsback & Zisserman, 2008), Oxford Pets (Parkhi et al., 2012), and Caltech101 (Fei-Fei et al., 2004). We reproduce baseline predictive models (Original) that are not exposed to synthetic data, with comparable performance (within $1\%$) of the baselines used in (Du et al., 2024; Zhang et al., 2023b). [^1] Baseline model checkpoints were not available so we reproduce them using the reported architecture and training recipe. We iteratively perform additional fine-tuning with synthetic training data that is equal in quantity across all conditions, use the same augmentations (random rotation, reflection, and center cropping), and report generalization performance as top-1 accuracy. Results are averaged over three runs. Additional experiment details are in Supplement A.1 and dataset details are in Supplement A.3. We fix hyperparameters (LTG weight, number of Epistemic Heads) across all datasets according to Supplement A.8.
@@
In Table 1, we compare to baselines under the Dataset Expansion task as defined in GIF (Zhang et al., 2023b). For parity with the architecture and quantity of synthetic data used by GIF, we fine-tune ResNet-50 models for 100 epochs and generate synthetic data equivalent to 20×20$\times$20 × the original dataset size for Caltech, Cars and Flowers, and 30×30$\times$30 × the original dataset size for Pets. GIF generates all synthetic data at once whereas we evenly distribute data generation throughout training epochs to better capture evolving model-based longtail challenges, see Supplement A.1 for details. In each expansion, synthetic examples per class are equal in quantity to the original number of real examples.
@@
In Table 2, we compare to baselines under the in-distribution synthetic data generation task as defined in Dream the Impossible (Dream-ID) (Du et al., 2024) and discussed in Supplement A.1. For parity with the architecture and quantity of synthetic data used by Dream-ID, we train ResNet-34 models for 100 epochs and generate 1k synthetic data per class. Models are trained on ImageNet-100 and evaluated on the same ImageNet-100, ImageNet-A, and ImageNet-V2 subsets as defined in Dream-ID and restated in Supplement A.1. In each benchmark, Longtail Guidance provides substantial improvements over existing synthetic data generation methods and data augmentation baselines. In particular, it provides an average $5.6$ points of additional top-1 accuracy compared to GIF or Dream-ID, and an average $15.0$ points of additional top-1 accuracy compared to text-prompted Stable Diffusion. Aggregate improvements from the Epistemic Head over alternative model-based longtail signals (energy, entropy) are mild but significant and come at negligible inference or training time cost as described in Supplement A.5. In Supplement A.4, we show that predictive model generalization can be further improved by engaging in more cycles of fine-tuning and synthetic data generation.
Strikingly, we find that predictive model generalization when fine-tuned on LTG-generated synthetic data eclipses generalization when fine-tuned on GIF-generated data at substantially lower synthetic data volumes: 20% for Caltech, 25% for Cars, 30% for Flowers, and 23% for Pets, for an overall average of 24.5%. Thus, LTG-generated data not only provide higher overall generalization improvements, but they do so at more than 4×4$\times$4 × the synthetic data efficiency. Furthermore, LTG-generated data provide value in both high synthetic data volume regimes (2030×20-30$\times$20 - 30 × original dataset size for GIF comparison) and low synthetic data volume regimes (less than 1×1$\times$1 × original dataset size for Dream-ID and ImageNet-LT comparisons).
@@
4.2 LTG Improves Longtail Performance
In Table 3, we train SOTA ViT-based models (LiVT) from scratch on ImageNet-LT according to the longtail-compensation approach of (Xu et al., 2023). In summary, training includes MAE pretraining followed by 100 epochs of BCE loss with a logit adjustment to account for class imbalance. Initial LiVT training is on real data only and uses many data augmentations including RandAugment, Mixup, and CutMix. Our baseline predictive model, LiVT, nearly matches the generalization performance (top-1 accuracy) of what is reported in (Xu et al., 2023) (LiVT Reproduced: $60.6$ vs LiVT Reported: $60.9$). We then fine-tune LiVT on $24k$ additional synthetic data evenly distributed across all $1k$ classes for $100$ epochs with $41e-$ 4 learning rate. We use the same Stable Diffusion 1.4 baseline and sampling details as in Section 4.1 (LiVT SD), and also compare Longtail Guidance with three types of model-based longtail signals (LiVT LTG {Energy, Entropy, Epistemic}). Overall generalization improvements on the balanced validation set are mild but significant ($+1.3$ top-1 accuracy). However, performance for longtail classes (Few) improves by four points for baseline diffusion ($+10\%$) but a marked $10$ points for Longtail Guidance ($+24\%$)!
Table 3: Top-1 Accuracy on ImageNet-LT by dataset split (Many, Med, Few) and overall (Acc). The validation set is balanced between all 1000 classes but training is highly imbalanced: Few classes have $5-19$ examples, Medium classes have $20-99$ examples, Many classes have $100-1200$ examples.